# Bongard-RWR+: Real-World Representations of Fine-Grained Concepts in Bongard Problems

### v1

@@ Appendix A Dataset variants
We also constructed Bongard-RWR+/LP, a dataset variant in which the image removal step is disabled (i.e., the flag $F$ in Algorithm 2 is set to false). Without this removal step, the greedy subset selection procedure for building $GS_{k}^{*}$ often draws from a smaller set of unique images within $GS_{k}$. In effect, matrices in Bongard-RWR+/LP exhibit lower average intra-side embedding similarity but are composed of fewer distinct images. The increased visual diversity within each matrix can facilitate recognition by presenting concepts in a broader range of views. However, because the same images are reused more frequently across matrices, there is overall less image-level diversity compared to Bongard-RWR+. We constructed dataset variants with $P=2,\ldots,6$, denoted Bongard-RWR+/L2, Bongard-RWR+/L3, etc. Example BPs from Bongard-RWR+/LP are shown in Figs. 15 – 19.
@@
Figure 11: Functional equivalence of generated images and real-world images. Performance trends in the I1S task on Bongard-RWR and Bongard-RWR+/LP for $P=2,3,4$.
Appendix B Extended results
B.1 Is image color necessary for concept recognition?
@@
B.2 Are generated images as effective as real ones?
We evaluated model performance on Bongard-RWR and Bongard-RWR+/LP in the I1S task for $P=2,3,4$. As shown in Fig. 11, results are comparable across both datasets, regardless of whether the model processes real-world or generated images. This finding is consistent with the experiment on the CS task presented in the main paper and reinforces our claim that generated images are equally effective as real-world images for probing visual reasoning capabilities of VLMs.
Table 4: Concept Selection on Bongard-RWR+/LP for $K=16$.
@@ B.2 Are generated images as effective as real ones?
B.3 Do models learn from demonstrations?
@@
Table 9: Concept Generation: Image-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from images in Bongard-RWR+ (using MiniCPM-o 2.6 $8\text{B}$) and Bongard-OpenWorld (using GPT-4V (wu2024bongardopenworld, , Appendix E, Table 4)).
@@
Table 10: Concept Generation: Text-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from image captions in Bongard-RWR+ (using Qwen2-VL $72\text{B}$ for prediction and InternVL2.5 $78\text{B}$ for captioning) and Bongard-OpenWorld (using ChatGPT for prediction and BLIP-2 w/ Fine-tuning for captioning (wu2024bongardopenworld, , Appendix E, Table 4)).
@@ B.4 Additional baselines
We evaluated the performance of supervised learning methods on the I1S and D1S tasks. The models operated on embeddings produced by pre-trained models. For I1S, raw images were encoded using ViT-L/14 dosovitskiy2021an , while for D1S, image captions produced by InternVL2.5 were embedded using fine-tuned MiniLM minilm2025hugging ; wang2020minilm . We considered an MLP with a single hidden layer applied to concatenated embeddings, Wild Relation Network (WReN) barrett2018measuring , which processes embedding pairs and aggregates them via summation, and SNAIL mishra2018a , an attention-based meta-learner. As shown in Table 11, all models performed near the random guess level on both Bongard-RWR+/TVT and Bongard-RWR+/TVT-Large, demonstrating that the proposed BPs present a significant challenge not only for VLMs but also for standard supervised learning approaches.
B.5 Performance on Concept Generation task
@@
Tables 5 – 8 evaluate model performance on the CG task, where models are prompted to produce free-form descriptions of the underlying concept for each matrix. Inputs were either raw images or image captions produced by InternVL2.5. Experiments were conducted on both Bongard-RWR+ and Bongard-RWR+/L6 datasets. We assess generated descriptions using standard NLP metrics: BLEU ($\text{BLEU}_{1}$ and $\text{BLEU}_{2}$papineni2002bleu , METEOR banerjee2005meteor , $\text{ROUGE}_{L}$ lin2004rouge , CIDEr lin2004rouge , and BERTScore (precision $P_{\text{BERT}}$, recall $R_{\text{BERT}}$, F1 $F_{\text{BERT}}$zhang2020bertscore . For BERTScore, we apply baseline rescaling [^1] https://github.com/Tiiiger/bert_score/blob/master/journal/rescale_baseline.md and the DeBERTa [^2] https://huggingface.co/microsoft/deberta-xlarge-mnli model he2021deberta , which has strong alignment with human evaluations. Higher scores indicate better performance across all metrics.
@@
MiniCPM-o 2.6 consistently performs best in image-based setups (Tables 5 – 6), ranking first in $5$ of $8$ metrics on Bongard-RWR+ and $6$ of $8$ on Bongard-RWR+/L6, though differences with other models are small. In text-based setups (Tables 7 – 8), Qwen2-VL leads in $5$ of $8$ metrics on Bongard-RWR+ and $4$ of $8$ on Bongard-RWR+/L6, again with relatively small gaps to competing models. We further compare top-performing models on Bongard-RWR+ with the best results reported for Bongard-OpenWorld (wu2024bongardopenworld, , Appendix E, Table 4), as shown in Tables 9 and 10. Regardless of input representation, models achieve significantly higher scores on Bongard-OpenWorld than on Bongard-RWR+, indicating that the concepts in our dataset are more challenging to recognize even for state-of-the-art models.
@@
B.6 Assessing difficulty of Bongard-RWR+ variants
@@
B.7 Qualitative analysis
Tables 13 and 14 present the top $5$ correctly recognized concepts per model for the I1S and D1S tasks, resp. In the image-based setup, we observe a diverse distribution of top concepts across the models, with little overlap. This variation suggests that each model relies on distinct reasoning strategies when processing visual input. For example, InternVL2.5 shows some capacity for identifying rules involving contour, similarity, and shape, whereas Qwen2-VL performs better on concepts related to count and angle. In contrast, the text-based setup yields more consistent results across models. Counting-related concepts dominate the top-performing examples, followed by shape and size.
@@ B.7 Qualitative analysis
Appendix C Limitations
@@
Our main experiments evaluated the ability of VLMs to solve BPs, alongside selected supervised learning baselines described in Appendix B. The results highlight clear limitations in the current models’ capacity for abstract visual reasoning, especially when dealing with fine-grained concepts, raising the need for more sophisticated approaches. In particular, multimodal reasoning models that integrate visual and textual information may offer a promising direction, analogous to how reasoning models like DeepSeek-R1 operate in the text domain. We envision the introduced datasets as valuable benchmarks for tracking progress in AVR capabilities.

### v2

@@ Appendix B Dataset Variants
We also constructed Bongard-RWR+/LP, a dataset variant in which the image removal step is disabled (i.e., the flag $F$ in Algorithm 2 is set to false). Without this removal step, the greedy subset selection procedure for building $GS_{k}^{*}$ often draws from a smaller set of unique images within $GS_{k}$. In effect, matrices in Bongard-RWR+/LP exhibit lower average intra-side embedding similarity but are composed of fewer distinct images. The increased visual diversity within each matrix can facilitate recognition by presenting concepts in a broader range of views. However, because the same images are reused more frequently across matrices, there is overall less image-level diversity compared to Bongard-RWR+. We constructed dataset variants with $P=2,\ldots,6$, denoted Bongard-RWR+/L2, Bongard-RWR+/L3, etc. Example BPs from Bongard-RWR+/LP are shown in Figs. 24 – 28.
@@
Figure 12: Functional equivalence of generated images and real-world images. Performance trends in the I1S task on Bongard-RWR and Bongard-RWR+/LP for a varying number of images per side $P=2,3,4$.
Appendix C Extended Results
C.1 Is image color necessary for concept recognition?
@@
Table 4: Performance correlation on generated and real-world images. Pearson correlation ($r$), coefficient of determination ($R^{2}$), regression parameters, mean absolute difference (MAD) and RMSE across $K\in\{2,4,8,16\}$ on Bongard-RWR and Bongard-RWR+/LP for $P\in\{2,3\}$.
Model $P$ $r$ $R^{2}$ Slope Intercept MAD RMSE
MiniCPM-o 2.6 8B 22 $0.997$ $0.994$ $0.971$ 0.005-0.005 $1.75$ $1.38$
MiniCPM-o 2.6 8B 33 $0.999$ $0.998$ $0.902$ $0.030$ $2.25$ $0.80$
InternVL2.5 78B 22 $0.994$ $0.987$ $1.204$ 0.190-0.190 $5.00$ $2.04$
InternVL2.5 78B 33 $0.993$ $0.987$ $1.419$ 0.377-0.377 $7.00$ $2.09$
Qwen2-VL 72B 22 $1.000$ $0.999$ $1.005$ 0.015-0.015 $1.25$ $0.43$
Qwen2-VL 72B 33 $1.000$ $0.999$ $1.004$ 0.035-0.035 $3.25$ $0.43$
LLaVA-Next 110B 22 $0.994$ $0.987$ $0.867$ $0.054$ $3.00$ $1.98$
LLaVA-Next 110B 33 $1.000$ $1.000$ $0.936$ $0.040$ $1.50$ $0.26$
C.2 Are generated images as effective as real ones?
To quantify the relationship presented in the paper (Fig. 9) more rigorously, we conducted a statistical analysis across all four models for both counts of panels per side ($P=2,3$) and all difficulty levels ($K=2,4,8,16$). We computed Pearson correlations and fitted linear regression models to assess how well performance on the generated-image dataset (Bongard-RWR+/LP) predicts performance on the real-image dataset (Bongard-RWR). The results are summarized in Table 4.
All models exhibit strong correlations between real-image and generated-image dataset performance ($r>0.99$), with $R^{2}$ consistently above $0.987$. For example, InternVL2.5 78B for $P=2$ shows $r=0.994$ and a regression slope of $1.204$, meaning that, e.g., a $10$ p.p. gain on Bongard-RWR+/L2 corresponds to roughly a $12$ p.p. gain on Bongard-RWR. Error measures including mean absolute difference (MAD $\leq 7.0$) and RMSE ($\leq 2.09$) confirm that Bongard-RWR+/LP is a slightly easier set, on average, but the relationship is stable and predictable across difficulty levels.
These results demonstrate that the generated-image dataset preserves the underlying difficulty of the real-image dataset. Consequently, performance improvements on Bongard-RWR+ can be expected to translate reliably to improvements on real-image BPs, supporting the validity of using generated images for AVR evaluation.
Furthermore, we evaluated model performance on Bongard-RWR and Bongard-RWR+/LP in the I1S task for a varying number of images per side $P=2,3,4$. As shown in Fig. 12, results are comparable across both datasets, regardless of whether the model processes real-world or generated images.
To further analyze the types of errors models make on real vs. generated images, we compared per-concept performance on Bongard-RWR (real) and Bongard-RWR+/LP for $P=3$. Across the $54$ concepts present in both datasets, performance was largely similar: in $40$ concepts ($74\%$) the difference was $\leq 10$ p.p., in $7$ concepts ($13\%$) differences were within 102010-20 p.p., and for $7$ concepts ($13\%$) differences exceeded $20$ p.p. Notably, in all larger-difference cases, models performed better on average on Bongard-RWR+/LP than on Bongard-RWR. In the small-difference cases, models performed better on Bongard-RWR for $5$ concepts and on Bongard-RWR+/LP for $2$ concepts. Overall, the results indicate that despite certain variation, the two datasets present a comparable level of difficulty.
These findings are consistent with the experiment on the CS task presented in the main paper and reinforces our claim that generated images are equally effective as real-world images for probing visual reasoning capabilities of VLMs.
Table 5: Concept Selection on Bongard-RWR+/LP for $K=16$ concepts.
@@ C.2 Are generated images as effective as real ones?
C.3 Do models learn from demonstrations?
@@
Table 10: Concept Generation: Image-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from images in Bongard-RWR+ (using MiniCPM-o 2.6 $8\text{B}$) and Bongard-OpenWorld (using GPT-4V (Wu et al., 2024, Appendix E, Table 4)).
@@
Table 11: Concept Generation: Text-based on Bongard-RWR+ vs. Bongard-OpenWorld. Comparison of concept descriptions generated from image captions in Bongard-RWR+ (using Qwen2-VL $72\text{B}$ for prediction and InternVL2.5 $78\text{B}$ for captioning) and Bongard-OpenWorld (using ChatGPT for prediction and BLIP-2 w/ Fine-tuning for captioning (Wu et al., 2024, Appendix E, Table 4)).
@@ C.4 Additional baselines
We evaluated the performance of supervised learning methods on the I1S and D1S tasks. The models operated on embeddings produced by pre-trained models. For I1S, raw images were encoded using ViT-L/14 (Dosovitskiy et al., 2021), while for D1S, image captions produced by InternVL2.5 were embedded using fine-tuned MiniLM (HuggingFace, 2021; Wang et al., 2020a). We considered an MLP with a single hidden layer applied to concatenated embeddings, Wild Relation Network (WReN) (Barrett et al., 2018), which processes embedding pairs and aggregates them via summation, and SNAIL (Mishra et al., 2018), an attention-based meta-learner. As shown in Table 12, all models performed near the random guess level on both Bongard-RWR+/TVT and Bongard-RWR+/TVT-Large, demonstrating that the proposed BPs present a significant challenge not only for VLMs but also for standard supervised learning approaches.
Figure 13: Model performance on Bongard-RWR+/L2 and the corresponding gain over Bongard-RWR+/L2-Impure in the CS task for $K\in\{2,4,8,16\}$. Substituting a human-approved image with one rejected during filtering hinders concept recognition.
C.5 Noisy images hinder concept recognition
Fig. 13 compares model performance between Bongard-RWR+/L2 and Bongard-RWR+/L2-Impure in the CS task across $K$ levels. Substituting an accepted image with a rejected one consistently reduces performance across models, with only a single case of comparable results (LLaVA-Next, $K=8$). The effect is strongest for InternVL2.5 (mean drop of $8.6$ p.p.) and Qwen2-VL ($5.7$ p.p.), moderate for LLaVA-Next ($4.6$ p. p.), and smallest for MiniCPM-o 2.6 ($3.9$ $p.$).
C.6 Performance on Concept Generation task
@@
Tables 6 – 9 evaluate model performance on the CG task, where models are prompted to produce free-form descriptions of the underlying concept for each matrix. Inputs were either raw images or image captions produced by InternVL2.5. Experiments were conducted on both Bongard-RWR+ and Bongard-RWR+/L6 datasets. We assess generated descriptions using standard NLP metrics: BLEU ($\text{BLEU}_{1}$ and $\text{BLEU}_{2}$(Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), $\text{ROUGE}_{L}$ (Lin, 2004), CIDEr (Lin, 2004), and BERTScore (precision $P_{\text{BERT}}$, recall $R_{\text{BERT}}$, F1 $F_{\text{BERT}}$(Zhang* et al., 2020). For BERTScore, we apply baseline rescaling [^4] https://github.com/Tiiiger/bert_score/blob/master/journal/rescale_baseline.md and the DeBERTa [^5] https://huggingface.co/microsoft/deberta-xlarge-mnli model (He et al., 2021), which has strong alignment with human evaluations. Higher scores indicate better performance across all metrics.
@@
MiniCPM-o 2.6 consistently performs best in image-based setups (Tables 6 – 7), ranking first in $5$ of $8$ metrics on Bongard-RWR+ and $6$ of $8$ on Bongard-RWR+/L6, though differences with other models are small. In text-based setups (Tables 8 – 9), Qwen2-VL leads in $5$ of $8$ metrics on Bongard-RWR+ and $4$ of $8$ on Bongard-RWR+/L6, again with relatively small gaps to competing models. We further compare top-performing models on Bongard-RWR+ with the best results reported for Bongard-OpenWorld (Wu et al., 2024, Appendix E, Table 4), as shown in Tables 10 and 11. Regardless of input representation, models achieve significantly higher scores on Bongard-OpenWorld than on Bongard-RWR+, indicating that the concepts in our dataset are more challenging to recognize even for state-of-the-art models.
@@
C.7 Assessing difficulty of Bongard-RWR+ variants
@@
C.8 Performance of the Similarity Classifier
The Similarity Classifier achieves high results in I1S experiments, as it exploits a fundamental property of BPs: a defining concept must hold for all images on a given side. This motivates the use of a max-aggregation heuristic, which assigns the test input to the side for which the maximum distance between individual image embeddings and the test image embedding (most outlying pair) is smaller. Fig. 15 illustrates the strength of the underlying design concept of SC.
Figure 15: An I1S example illustrating the strength of the Similarity Classifier. Left: Polygons. Right: Curvilinear figures. The right side shows a test image, and the left side presents a plot of distances from the test image to images on both sides in the ViT-L/14 embedding space. The ice sculpture is farther from the brick wall in the Left side than from any image in the Right side, leading to a higher probability of being classified as a curvilinear figure. In contrast, InternVL2.5 relies on local features and doesn’t take into account all side images, resulting in an incorrect classification: “The test image shows a symmetric ice sculpture, matching the symmetric patterns seen in the LEFT class images. Conversely, the RIGHT class images feature more natural, organic, or architectural elements that lack symmetry.”
(a) Image-to-Side
(b) Description-to-Side
Figure 16: Per-concept difficulty distributions. Average number of correctly solved instances (out of $200$) per concept across all models in Bongard-RWR+ for (a) I1S and (b) D1S tasks.
@@ C.8 Performance of the Similarity Classifier
Gemini 2.5 Pro $1.00$ Count One figure Two figures
$0.93$ Size Large figures Small figures
$0.93$ Shape Triangles Circles
$0.92$ Angle A sharp projection No sharp projection
$0.91$ Similarity All figures of the same color Figures of different colors
Claude Sonnet 4.5 $0.97$ Count One figure Two figures
$0.92$ Similarity All figures of the same color Figures of different colors
$0.91$ Angle A sharp projection No sharp projection
$0.90$ Shape Triangles Circles
$0.88$ Count Three parts Four parts
GPT 5.1 $1.00$ Count One figure Two figures
$0.98$ Position Triangle inside of the circle Circle inside of the triangle
$0.95$ Shape Triangles Circles
$0.94$ Angle A sharp projection No sharp projection
$0.93$ Similarity All figures of the same color Figures of different colors
@@
Gemini 2.5 Pro 1.00 Count One figure Two figures
1.00 Count Three parts Four parts
0.98 Position Triangle inside of the circle Circle inside of the triangle
0.97 Similarity All figures of the same color Figures of different colors
0.96 Size Triangle larger than circle Triangle smaller than circle
Claude Sonnet 4.5 0.99 Count One figure Two figures
0.98 Similarity All figures of the same color Figures of different colors
0.94 Shape Polygons Curvilinear figures
0.90 Count Three parts Four parts
0.88 Contour Large total line length Small total line length
GPT-5.1 1.00 Count One figure Two figures
0.98 Position Triangle inside of the circle Circle inside of the triangle
0.96 Similarity All figures of the same color Figures of different colors
0.95 Shape Polygons Curvilinear figures
0.94 Count Three parts Four parts
C.9 Qualitative analysis
To understand how concept difficulty varies across input modalities, we analyzed the average number of correctly solved instances per concept for all models. Fig. 16 presents the distribution of concept difficulty for both I1S and D1S tasks.
For I1S, each concept was solved in 7979109109 instances (out of $200$) on average. This relatively narrow range indicates that most concepts remain uniformly challenging for models in the visual domain, which is consistent with their near-random performance levels. One noteworthy case is “Shading thicker on the right side” (concept #63\#63), which appears among top-5 solved concepts in I1S ($109$) but is relatively difficult in D1S ($92$), highlighting the relevance of visual input in identifying fine-grained concepts.
In contrast, D1S exhibits a much broader difficulty range, from $70$ to $182$ solved instances, reflecting high performance on certain concepts that appear to be easier. Notably, the top-5 solved concepts include counting-based ones such as “three parts” or “one figure”, which may be easier to capture in the text domain.
Tables 14 and 15 present the top $5$ correctly recognized concepts per model for the I1S and D1S tasks, resp. In the image-based setup, we observe a diverse distribution of top concepts across the models, with little overlap. This variation suggests that each model relies on distinct reasoning strategies when processing visual input. For example, InternVL2.5 shows some capacity for identifying rules involving contour, similarity, and shape, whereas Qwen2-VL performs better on concepts related to count and angle. In contrast, the text-based setup yields more consistent results across models. Counting-related concepts dominate the top-performing examples, followed by shape and size.
@@ C.9 Qualitative analysis
(a) Outline vs. solid figures
(b) Large vs. small total line length
Figure 17: Failure mode examples. (a) InternVL2.5’s prediction: “Convex hull of figure elongated vs. Convex hull of figure compact”; explanation: “The images on the left side feature objects or shapes that are more elongated, such as the person’s face, the mirror, the stones, and the clock. […]”. Assessment: The omission of the pie in the description of the left side illustrates error type (1). Furthermore, while faces and mirrors indeed appeared on the left side, the stones and clock were on the right, exemplifying error type (2). (b) Qwen2-VL’s prediction: “Closed lines vs. Open lines”; explanation: “The left side of the image contains natural landscapes […]”. Assessment: Although some left-side images indeed contain landscapes, landscapes also appeared on the right side, exemplifying error type (3). The explanation also reflects type (4) error, as it relies on surface-level features while ignoring the underlying geometric property.
C.10 Sources of error and failure modes
In Section 4 (Concept Selection) we noted that concepts such as Contour, Rotation, and Angle pose particular difficulties because they rely on subtle geometric cues or precise spatial relations that current models struggle to capture reliably. To further explore these weaknesses, we deconstructed the concept-group aggregates (Fig. 6) and reviewed model performance on specific concepts in the CS task with $K=4$. For example, InternVL2.5 achieved consistently strong results on all concepts within the Size, Shape, and Branching groups, but struggled with certain concepts from the three more difficult groups mentioned above (Contour, Rotation, and Angle):
  1. 1.
    For Angle ($69\%$ average accuracy), the model performed particularly poorly on concept #76\#76 “Long sides concave vs. Long sides convex” ($16\%$), which requires abstracting beyond surface-level features to recognize a geometric property across diverse real-world scenes.
  2. 2.
    For Rotation ($73\%$), performance dropped on concept #17\#17 “An acute angle directed inward vs. No acute angle directed inward” ($43\%$), again pointing to difficulties in detecting precise spatial arrangements.
  3. 3.
    For Contour ($56\%$), concept #3\#3 “Outline figures vs. Solid figures” ($22\%$) was especially challenging, suggesting difficulties in recognizing fine-grained details such as object boundaries.
Solving these tasks requires abstraction beyond immediate object features – precisely the type of reasoning our benchmark aims to measure.
We also investigated typical model error patterns in the CS task by manually reviewing prediction explanations. Several recurring failure modes emerged: (1) predicting a concept that applies only to a subset of images on a given side, (2) confusing the sides images belong to (3) predicting a concept for one side while ignoring that it also applies to the opposite side, (4) relying on shallow surface-level features rather than deep image understanding required to detect fine-grained concepts. Fig. 17 presents the examples.
Model CS D1S I1S
Gemini 2.5 Pro 0.82 0.70 0.65
GPT 5.1 $0.81$ $0.67$ $0.65$
Claude Sonnet 4.5 $0.76$ $0.67$ $0.63$
Figure 18: Proprietary model results on Bongard-RWR+ in the CS ($K=16$), D1S, and I1S tasks.
C.11 Frontier proprietary models
To contextualize the performance of open VLMs, we conducted a limited evaluation of three state-of-the-art proprietary models—Claude Sonnet 4.5, Gemini 2.5 Pro, and GPT 5.1—on Bongard-RWR+ across the I1S, D1S, and CS ($K=16$) tasks. As shown in Fig. 18(a), these models consistently outperform open ones (cf. Fig. 5 and Table 2), with particularly strong and stable performance on the Shape, Size, and Count concept groups (Fig. 18(b)). Their advantage is most pronounced in the vision-intensive I1S task, where all open models fall to near-chance accuracy despite performing reasonably well on the same groups in CS and, to a lesser extent, D1S. Nevertheless, even the best-performing proprietary model (Gemini 2.5 Pro) reaches only $65\%$ on I1S, $70\%$ on D1S, and $82\%$ on CS, indicating that Bongard-RWR+ presents a significant challenge also for frontier models.
Appendix D Limitations and future work
@@
Our main experiments evaluated the ability of VLMs to solve BPs, alongside selected supervised learning baselines described in Appendix C.4. The results highlight clear limitations in the current models’ capacity for abstract visual reasoning, especially when dealing with fine-grained concepts, raising the need for more sophisticated approaches. In particular, multimodal reasoning models that integrate visual and textual information may offer a promising direction, analogous to how reasoning models like DeepSeek-R1 operate in the text domain. We envision the introduced datasets as valuable benchmarks for tracking progress in AVR capabilities and, more broadly, in multi-image reasoning.